# Resonant spectral cascade in pulsatile flow triggered by arterial geometry

### v1

@@ Abstract
The increasing geometric complexity of arteries with age is a well-established risk factor for cardiovascular disease, yet the precise physical mechanisms by which geometry interacts with blood flow remain debated. While complex geometry is known to increase viscous resistance, its active role in modulating flow dynamics is not fully understood. Here we solve a mathematical model to show that arterial geometry can trigger a resonant transfer of energy to short-wavelength components of the flow. The investigation, conducted over a physiological range of Womersley numbers (Wo, a dimensionless measure of pulsation frequency), reveal a dual dynamic. The global wave energy consistently decays, confirmed by a negative growth rate (G ¡ 0), indicating that the flow does not become exponentially unstable. However, a spectral broadening ratio (R), which quantifies the energy in high-wavenumber versus low-wavenumber modes, exhibits a sharp, non-monotonic peak at an intermediate Wo. This result identifies a resonant frequency at which geometry is maximally efficient at generating spectral complexity, even as the overall flow attenuates. These findings reframe the role of arterial geometry from a passive dissipator to an active modulator of the flow’s spectral content, suggesting that spectral diagnostics could provide a sensitive marker for vascular health.
@@ Introduction
Blood flow in large arteries exhibits complex dynamics resulting from the interaction of pulsatile forcing, viscous dissipation, elastic wall response, and vessel geometry. High-resolution computational fluid dynamics (CFD) and particle image velocimetry (PIV) experiments have demonstrated that these factors can lead to complex, turbulence-like flow states characterized by a spectral cascade of kinetic energy, even under physiologic conditions [1, 2]. Factors such as vessel curvature [3], wall compliance [5, 4], and pathologic features like stenosis [6] are known to fundamentally alter local shear distributions and energy transport.
Curved and torsional arteries are not passive conduits. Classical studies of pulsatile flow in curved pipes established that curvature generates Dean-type secondary vortices and periodic shifts in velocity profiles that reshape wall shear stress [7]. Later analyses of fully developed oscillatory flow quantified how curvature amplitude controls phase lag and energy loss [8]. Laboratory experiments confirmed that increasing Dean number intensifies cross-stream motion and introduces nonuniform damping along the wall [9]. More recently, canonical curved-cavity experiments demonstrated that the interaction between Reynolds, Dean, and Womersley numbers produces distinct vortex morphologies analogous to those in intracranial aneurysms [10]. Computational models of coronary flow further showed that geometric curvature directly influences fractional flow reserve and hemodynamic stability under pulsatile forcing [3, 11]. Collectively, these results highlight that geometry-induced dispersion and secondary flow development are central to the stability of arterial pulsations.
@@
In this sense, the study challenges the prevailing view that viscous dissipation universally stabilizes pulsatile flow. It suggests instead that under specific combinations of curvature, torsion, and pulsation frequency, geometry can act as a driver rather than a dissipator. While variable-coefficient KdV models have been foundational in fields like nonlinear optics and shallow-water wave theory, the present work originally applies this framework to pulsatile hemodynamics. Our approach unifies classical curved-pipe hydrodynamics [7, 8, 9], modern computational hemodynamics [3, 12, 13], and fractional modeling theory [18, 19, 20] into a single analytical–computational pipeline. By turning geometry from a passive constraint into an active variable, this work reframes arterial stability as a question of parametric energy exchange, opening a new path for interpreting instability in pulsatile biological flows.
@@ Methods
This section derives a nonlinear one-dimensional model for pulsatile blood flow in geometrically complex arteries from the Navier–Stokes equations, followed by the computational workflow used for stability and chaos diagnostics.
Governing Equations.
Let $\rho$ [kg m-3] and $\mu$ [Pa s] denote blood density and dynamic viscosity. In cylindrical coordinates $(r,\theta,z)$, the velocity components are $(v_{r},v_{\theta},v_{z})$ [m s-1] and the pressure field is $p$ [Pa]. The incompressible Navier–Stokes system reads
$\frac{1}{r}\frac{\partial}{\partial r}(rv_{r})+\frac{1}{r}\frac{\partial v_{\theta}}{\partial\theta}+\frac{\partial v_{z}}{\partial z}$ $=0,$ (1)
@@
For straight, axisymmetric vessels $(\partial/\partial\theta=0,v_{\theta}=0)$, Eq. (2) simplifies to
$\rho\left(\frac{\partial v_{z}}{\partial t}+v_{r}\frac{\partial v_{z}}{\partial r}+v_{z}\frac{\partial v_{z}}{\partial z}\right)=-\frac{\partial p}{\partial z}+\mu\!\left[\frac{1}{r}\frac{\partial}{\partial r}\!\left(r\frac{\partial v_{z}}{\partial r}\right)\right].$ (3)
Base Oscillatory (Womersley) Flow: Neglecting convective terms in Eq. (3) gives the linear unsteady Stokes equation. For a harmonic pressure gradient $\partial p/\partial z=\Re\{\hat{P}e^{i\omega t}\}$, the exact velocity solution is
@@
where $\mathrm{Wo}=R_{0}\sqrt{\omega/\nu}$ is the Womersley number, $\nu=\mu/\rho$, and $J_{0}$ is the Bessel function of the first kind. This oscillatory profile forms the base state for subsequent perturbation analysis.
Weakly Nonlinear Reduction: To capture finite-amplitude effects neglected by Eq. (4), we expand the axial velocity as
@@
where $x$ is the centerline coordinate, $\epsilon\ll 1$ is a small parameter, $A(x,t)$ is the slowly varying wave amplitude, and $\phi(r)$ is the normalized radial eigenfunction of the linear operator from Eq. (3). Substituting Eq. (5) into Eq. (3), projecting onto $\phi(r)$, and retaining $\mathcal{O}(\epsilon^{2})$ terms yields the amplitude evolution equation
@@
The nonlinear term $C(x)A(\partial A/\partial x)$ arises directly from the convective acceleration $v_{z}(\partial v_{z}/\partial z)$ in Eq. (3) and represents amplitude-dependent wave steepening. The dispersive coefficient $D(x)$ comes from higher-order viscous and geometric corrections, and $K(x)\mathcal{D}^{\gamma(x)}A$ models distributed geometric damping through the Riesz fractional derivative
$\mathcal{F}[\mathcal{D}^{\gamma(x)}A]=|k|^{1+\gamma(x)}\hat{A}(k),$ (7)
where $\mathcal{F}$ denotes the Fourier transform.
Dimensionless Form and Geometric Modulation.
Introducing scales $L_{0}$, $T_{0}=1/\omega$, $A_{0}$, and dimensionless variables $\xi=x/L_{0}$, $\tau=t/T_{0}$, $a_{1}=A/A_{0}$, gives
@@
with $\alpha_{0}$, $\beta(\xi)$, and $\eta(\xi)$ representing nonlinearity, dispersion, and damping. The general form includes a Riesz fractional operator, but for this study, the order was fixed to $\gamma_{0}=0$, which reduces the operator to a form proportional to the standard viscous term. The nonlinearity was held constant ($\alpha_{0}=1$). Periodic spatial modulations mimic curvature or torsion:
$\beta(\xi)=\beta_{0}[1+\varepsilon_{\beta}\cos(q\xi)],\quad\eta(\xi)=\eta_{0}[1+\varepsilon_{\eta}\cos(q\xi)],$
where $q$ is the geometric wavenumber and $\varepsilon_{i}$ are modulation amplitudes $(0\!\leq\!\varepsilon_{i}\!\leq\!0.3)$. The baseline scalings are physically motivated: $\beta_{0}\!\propto\!\mathrm{Wo}^{-2}$ arises from the balance of inertial and viscous effects in the underlying Womersley flow, while $\eta_{0}=\eta_{\text{ref}}(1+0.1/\mathrm{Wo})$ models frequency-dependent dissipation from viscous boundary layers, with a reference damping value of $\eta_{\text{ref}}=0.005$.
Initial and Boundary Conditions: The system is solved on a periodic domain of length $L_{g}$. The initial condition is a physiologically informed multi-harmonic waveform given by $a_{1}(\xi,0)=\sum_{n=1}^{3}A_{n}\sin(nk_{0}\xi+\phi_{n})$ with amplitude ratios $A_{2}/A_{1}=0.3$ and $A_{3}/A_{1}=0.1$. For the detailed representative case shown in Figures 2-6, a fundamental wavenumber of $k_{0}=0.5$ was used. For the parametric sweep summarized in Figures 7-9, $k_{0}=1.0$ was used. The use of a Fourier-based solver inherently enforces periodic boundary conditions for all simulations.
Parameter Range: Simulations span $\mathrm{Wo}\in\{2,5,10,15,20\}$ and modulation amplitudes $\varepsilon_{i}\leq 0.3$. The dimensionless geometric period was held constant at $qL_{g}=4\pi$. This range covers small- to large-artery regimes [15] and conditions for curvature-driven resonance [16, 17].
Numerical Integration Scheme: Equation (8) is solved numerically using a split-step Fourier pseudospectral method. The domain of length $L_{g}$ is discretized into $N$ uniformly spaced points. The nonlinear term, $\alpha(\xi)a_{1}\partial_{\xi}a_{1}$, is calculated in real space and advanced in time using a fourth-order Runge-Kutta scheme. The linear part, containing the dispersive and fractional damping terms, is handled in Fourier space.
For computational efficiency, the spatially-varying linear operator is approximated by using the spatial mean of the coefficients, $\bar{\beta}=\langle\beta(\xi)\rangle$ and $\bar{\eta}=\langle\eta(\xi)\rangle$. This defines a constant-coefficient linear operator in Fourier space, $L(k)=-i\bar{\beta}k^{3}-\bar{\eta}|k|^{1+\gamma_{0}}$. The solution is advanced each timestep by applying an exact exponential propagator based on this mean operator, $\hat{a}_{1}(k,\tau+\Delta\tau)=e^{\Delta\tau L(k)}\hat{a}_{1}(k,\tau)$. This approach robustly captures the average dispersive and damping effects while significantly improving numerical performance.
@@
To quantify the system’s dynamics, we employ two primary diagnostics. First, the global energy of the wave is defined as $I_{2}(\tau)=\int_{0}^{L_{g}}a_{1}^{2}\,d\xi$. Its instantaneous logarithmic growth rate,
$G(\tau)=\frac{d}{d\tau}\ln I_{2}(\tau),$ (9)
is used to identify periods of energy growth ($G>0$) or decay ($G<0$), signaling global stability.
Second, to measure the transfer of energy to small scales, we define the spectral broadening ratio, $R(\tau)$. This is the ratio of energy in high-wavenumber modes to that in low-wavenumber modes:
$R(\tau)=\frac{E_{\text{high}}}{E_{\text{low}}}=\frac{\int_{k_{c}}^{k_{\mathrm{max}}}|\hat{a}_{1}(k,\tau)|^{2}dk}{\int_{0}^{k_{c}}|\hat{a}_{1}(k,\tau)|^{2}dk},$
where the cutoff wavenumber is defined relative to the initial condition’s fundamental wavenumber as $k_{c}=1.5\times(3k_{0})$, separating the initial energy-containing modes from the high-frequency cascade. A value of $R>1.5$ is taken as a threshold indicating a significant spectral cascade and the onset of a broadband, complex state.
The dimensionless amplitude $a_{1}$ corresponds to physical velocity perturbation $v_{z}^{\prime}(x,t)=\\ A_{0}a_{1}(x/L_{0},t/T_{0})$. In the limit $\varepsilon_{i}=0$ and $\eta=0$, Eq. (8) recovers the integrable constant-coefficient KdV equation, conserving invariants $I_{1}$ and $I_{3}$. The solver reproduces the analytical Womersley velocity field for vanishing nonlinearity, validating both physical and numerical consistency. All algorithms were implemented in Python using NumPy and FFTW; code and data are available for reproducibility.
Solution Approach
The numerical solution of the governing equation is performed using the split-step Fourier pseudospectral scheme detailed above, implemented in Python. The core computations and Fast Fourier Transforms rely on the NumPy and SciPy libraries, while the parameter sweep results were managed with pandas and all figures were generated using Matplotlib. All spatial derivatives are evaluated spectrally, which inherently enforces the periodic boundary conditions. Temporal integration continues until the diagnostic quantities reach a quasi-steady state.
The solver is verified by confirming that it reproduces analytical solutions for the Korteweg-de Vries equation in the constant-coefficient limit ($\varepsilon_{i}=0$) and conserves the energy invariant $I_{2}$ in the absence of damping ($\eta=0$). Numerical convergence was established by refining the spatial grid spacing $\Delta\xi$ and timestep $\Delta\tau$ until the relative change in the total energy $I_{2}$ between runs was less than $10^{-5}$. The scheme does not employ an explicit de-aliasing procedure; however, an additional check for numerical accuracy was performed by monitoring the energy in the highest-wavenumber modes, which remained a negligible fraction of the total energy, providing confidence that aliasing errors did not contaminate the physical cascade. All source code and data are available to ensure full reproducibility. The key numerical and physical parameters used in the simulations are summarized in Table 1.

### v2

# Resonant spectral cascade in Womersley flow triggered by arterial geometry

@@ Abstract
Age-related arterial remodeling is dominated by progressive loss of elastic-fiber function and concomitant stiffening, and in many vascular beds it is also accompanied by measurable geometric remodeling (e.g., elongation and tortuosity). These changes are clinically relevant because they modify pulsatile phase relationships, near-wall shear, and axial transport, yet the precise physical mechanisms by which geometry modulates spectral energy redistribution remain insufficiently resolved. While complex geometry is known to increase viscous resistance, its active role in modulating flow dynamics is not fully understood. Here we solve a mathematical model to show that arterial geometry can trigger a resonant transfer of energy to short-wavelength components of the flow. The investigation, conducted over a physiological range of Womersley numbers (Wo, a dimensionless measure of pulsation frequency), reveal a dual dynamic. The global wave energy consistently decays, confirmed by a negative growth rate (G ¡ 0), indicating that the flow does not become exponentially unstable. However, a spectral broadening ratio (R), which quantifies the energy in high-wavenumber versus low-wavenumber modes, exhibits a sharp, non-monotonic peak at an intermediate Wo. This result identifies a resonant frequency at which geometry is maximally efficient at generating spectral complexity, even as the overall flow attenuates. These findings reframe the role of arterial geometry from a passive dissipator to an active modulator of the flow’s spectral content, suggesting that spectral diagnostics could provide a sensitive marker for vascular health.
@@ Introduction
Blood flow in large conduit arteries arises from the interaction of pulsatile forcing, viscous dissipation, vessel geometry, and wall mechanics. Even under physiologic Reynolds numbers ($Re\approx O(10^{2})$), these interacting effects can generate flow states exhibiting nontrivial spectral structure beyond that predicted by classical steady pipe-flow theory. High-resolution computational fluid dynamics (CFD) and particle image velocimetry (PIV) studies have reported turbulence-like spectral features and departures from Kolmogorov scaling in physiologic arterial conditions [1, 2]. Geometric factors such as curvature [3], compliance [5, 4], and localized pathologies such as stenosis [6] are known to modify shear distribution, phase relationships, and energy transport. However, rather than attempting to reproduce fully resolved arterial turbulence, the present work focuses on a multi-harmonic pulsatile base state, following the Womersley flow shown by Saqr et al  [1] to exhibit sensetive-dependence on initial condistions and non-Kolmogorov scaling [7]. The present work investigates how geometry alone can induce transient spectral redistribution within this controlled framework. This mechanistic perspective isolates a minimal pathway by which pulsatile inertia and geometric modulation can influence axial spectral content in the moderate-to-high Womersley regime relevant to large arteries.
Age-associated arterial remodeling is multifactorial and should not be interpreted as geometric change alone. A primary and widely documented effect of aging is progressive stiffening driven by elastic-fiber degradation and cumulative damage to elastin and microfibrillar structure [8]. At the same time, multiple imaging and morphometric studies report that geometric remodeling such as elongation and tortuosity can increase with age in large vessels and selected vascular territories, although its prevalence and severity vary substantially across individuals and beds. For example, quantitative analysis of archived aortograms shows a significant age-correlation of abdominal aortic tortuosity [10], and early clinical hemodynamic discussions similarly noted a rising incidence of tortuosity beyond midlife and its potential implication for abnormal wall shear [11]. Age-stratified CT angiography of the vertebral artery further reports high tortuosity prevalence in older groups, consistent with age-related geometric remodeling in specific arterial segments [9]. In this context, the present study does not claim that aging universally increases geometric complexity; rather, it isolates geometry as an independent mechanistic pathway that can modulate pulsatile spectral redistribution when geometric remodeling is present, while recognizing that stiffness remains the dominant aging hallmark [8].
In reality, curved and torsional arteries are not passive conduits. Classical studies of pulsatile flow in curved pipes established that curvature generates Dean-type secondary vortices and periodic shifts in velocity profiles that reshape wall shear stress [12]. Later analyses of fully developed oscillatory flow quantified how curvature amplitude controls phase lag and energy loss [13]. Laboratory experiments confirmed that increasing Dean number intensifies cross-stream motion and introduces nonuniform damping along the wall [14]. More recently, canonical curved-cavity experiments demonstrated that the interaction between Reynolds, Dean, and Womersley numbers produces distinct vortex morphologies analogous to those in intracranial aneurysms [15]. Computational models of coronary flow further showed that geometric curvature directly influences fractional flow reserve and hemodynamic stability under pulsatile forcing [3, 16]. Collectively, these results highlight that geometry-induced dispersion and secondary flow development are central to the stability of arterial pulsations.
@@
In this sense, the presented work aims to refine the classical understanding of viscous stabilization in pulsatile flow. Viscous effects remain globally dissipative in the model, as reflected by the negative total energy growth rate $G(\tau)$ across the parameter ranges examined. However, the analysis aims to investigate geometry-modulated dispersion and if it can transiently redistribute spectral energy across axial wavenumbers before viscous damping dominates. Under specific combinations of curvature, torsion, and pulsation frequency, this redistribution could manifest as temporary amplification of higher-wavenumber content within an overall energy-decaying system. Geometry therefore does not replace viscous stabilization in the present work, but modulates the pathway by which energy is redistributed prior to eventual attenuation.
Variable-coefficient KdV-type formulations have long provided reduced descriptions of dispersive transport in nonlinear optics and shallow-water hydrodynamics. The present work adapts this framework to pulsatile hemodynamics, embedding curvature- and torsion-dependent proxy coefficients within a Womersley-scaled setting. In doing so, it connects classical curved-pipe hydrodynamics [12, 13, 14], contemporary computational hemodynamics [3, 17, 18], and recent fractional modeling approaches [23, 24, 25] within a unified analytical–computational structure. By treating geometry as a parametric modulator of dispersive and dissipative balance, the study reframes arterial spectral behavior as a problem of transient energy redistribution rather than classical linear instability, thereby offering a mechanistic perspective on pulsatile biological flows.
The present work is intended to represent large, conduit-type systemic arteries in the moderate-to-high Womersley regime, rather than small muscular arteries or the microcirculation. Specifically, the parameter range $\mathrm{Wo}=5$1515 corresponds to vessels such as the proximal aorta, carotid, and major femoral arteries under resting to moderately elevated heart rates. For a characteristic radius $R_{0}\sim 5$1515 mm, kinematic viscosity $\nu\approx 3\times 10^{-6}\,\mathrm{m}^{2}/\mathrm{s}$, and pulsation frequency $\omega=2\pi f$ with $f\sim 1$22 Hz, the resulting Womersley number $\mathrm{Wo}=R_{0}\sqrt{\omega/\nu}$ naturally lies in this interval, where $R_{0}$ denotes the reference pipe (or arterial) radius, $\omega$ the pulsation frequency, and $\nu$ the kinematic viscosity. The wall is treated as effectively rigid in the present formulation in order to isolate geometry-induced dispersive and dissipative modulation from compliance effects; this assumption is most appropriate for large arteries over short transient timescales where inertial and pulsatile effects dominate wall motion to leading order. Smaller arteries (with $\mathrm{Wo}<5$) and highly compliant vessels would require explicit fluid–structure interaction modelling, which is beyond the scope of the present work. The conclusions drawn herein should therefore be interpreted as applying to large systemic arteries in which pulsatile inertia and geometric modulation coexist within a moderate Womersley-number regime.
@@ Methods
This section summarizes (i) the reduced one-dimensional evolution model used to represent weakly nonlinear pulsatile dynamics in a conduit-type artery, (ii) the nondimensional parameterization linking the reduced coefficients to Womersley number, and (iii) the numerical scheme and diagnostics used to quantify global decay and transient spectral redistribution.
Governing equations and pulsatile base state
Let $\rho$ [kg m-3] and $\mu$ [Pa s] denote density and dynamic viscosity. In cylindrical coordinates $(r,\theta,z)$, the velocity is $\mathbf{v}=(v_{r},v_{\theta},v_{z})$ [m s-1] and the pressure is $p$ [Pa]. Incompressibility gives
$\frac{1}{r}\frac{\partial}{\partial r}(rv_{r})+\frac{1}{r}\frac{\partial v_{\theta}}{\partial\theta}+\frac{\partial v_{z}}{\partial z}=0.$ (1)
The axial momentum equation is
@@
For axisymmetric flow $(\partial/\partial\theta=0)$ this reduces to
$\rho\left(\frac{\partial v_{z}}{\partial t}+v_{r}\frac{\partial v_{z}}{\partial r}+v_{z}\frac{\partial v_{z}}{\partial z}\right)=-\frac{\partial p}{\partial z}+\mu\left[\frac{1}{r}\frac{\partial}{\partial r}\!\left(r\frac{\partial v_{z}}{\partial r}\right)+\frac{\partial^{2}v_{z}}{\partial z^{2}}\right].$ (3)
Here $z$ denotes the axial coordinate. In the reduced one-dimensional evolution equation below, the coordinate $x$ (and its nondimensional form $\xi=x/L_{0}$) is used for the same axial direction.
To define the laminar pulsatile reference state, we impose fully developed oscillatory pipe-flow assumptions: $v_{r}=0$, $v_{\theta}=0$, and $\partial v_{z}/\partial z=0$. Under these assumptions, Eq. (1) is satisfied and the convective terms vanish identically in Eq. (3), yielding the unsteady Stokes problem driven by a harmonic pressure gradient. For $\partial p/\partial z=\Re\{\hat{P}e^{i\omega t}\}$, the classical Womersley solution is
@@
where $\mathrm{Wo}=R_{0}\sqrt{\omega/\nu}$ is the Womersley number, $\nu=\mu/\rho$, $R_{0}$ is the reference radius, and $J_{0}$ is the Bessel function of the first kind. This oscillatory profile is taken as the laminar base state about which weakly nonlinear axial evolution is modeled.
Reduced one-dimensional evolution model
To represent weakly nonlinear axial evolution about the pulsatile base state (4), we expand
@@
where $\epsilon\ll 1$, $A(x,t)$ is a slowly varying amplitude, and $\phi(r)$ is a normalized radial eigenfunction associated with the linear operator in Eq. (3). Retaining $\mathcal{O}(\epsilon^{2})$ contributions and projecting onto $\phi(r)$ yields a KdV-type amplitude equation with dispersion and dissipation terms representing effective axial transport physics:
@@
The nonlinear steepening term is the reduced manifestation of convective acceleration in Eq. (3). Dissipation is represented through a Riesz fractional operator, defined in Fourier space by
$\mathcal{F}\!\left[\mathcal{D}^{\gamma(x)}A\right]=|k|^{1+\gamma(x)}\hat{A}(k),$ (7)
where $\hat{A}$ is the Fourier transform of $A$. The derivation of this variable-coefficient fractional KdV modelis obtained by expanding about the oscillatory Womersley base flow, introducing a weakly nonlinear long-wave perturbation, and projecting onto the leading radial eigenmode. The fractional dissipation term is used here as a reduced-order representation of multiscale nonlocal damping induced by geometric complexity; full details are given in Supplementary Information (A.1–A.10.).
Nondimensional form and coefficient parameterization
Introduce reference scales $L_{0}$, $T_{0}=1/\omega$, and $A_{0}$, and define
$\xi=\frac{x}{L_{0}},\qquad\tau=\frac{t}{T_{0}},\qquad a_{1}(\xi,\tau)=\frac{A(x,t)}{A_{0}}.$ (8)
The nondimensional evolution equation used throughout the manuscript is
@@
where $\alpha_{0}$ is the (constant) nonlinear coefficient, $\beta(\xi)$ is the dispersion coefficient, and $\eta(\xi)$ is the dissipation coefficient. In this study we fix $\gamma_{0}=0$, so the dissipative term has Fourier symbol $|k|^{1+\gamma_{0}}=|k|$ per Eq. (7). We also set $\alpha_{0}=1$.
To represent axial variability associated with geometric remodeling (curvature, tortuosity, and roughness) in a reduced-order manner, we parameterize $\beta(\xi)$ and $\eta(\xi)$ by bounded sinusoidal modulations
$\beta(\xi)=\beta_{0}\left[1+\varepsilon_{\beta}\cos(q\xi)\right],\qquad\eta(\xi)=\eta_{0}\left[1+\varepsilon_{\eta}\cos(q\xi)\right],$ (10)
with modulation amplitudes $0\leq\varepsilon_{\beta},\varepsilon_{\eta}\leq 0.3$ and geometric wavenumber $q$.
The baseline coefficients are scaled with Womersley number as
$\beta_{0}=\beta_{\mathrm{ref}}\,\mathrm{Wo}^{-2},$ (11)
and
$\eta_{0}=\eta_{\mathrm{ref}}\left(1+\frac{C_{\eta}}{\mathrm{Wo}}\right),$ (12)
where $\eta_{\mathrm{ref}}=0.005$ and $C_{\eta}=0.1$ are constants used in all simulations. The dimensional scaling leading to this $\mathrm{Wo}^{-2}$ dependence is derived explicitly in the Supplementary Information (Sec. A.10). The $\mathrm{Wo}^{-1}$ correction in Eq. (12) is an order-of-magnitude representation of the oscillatory Stokes boundary-layer scaling $\delta\sim\sqrt{2\nu/\omega}$, implying $\delta/R_{0}=O(\mathrm{Wo}^{-1})$ and therefore a weak increase of near-wall viscous contribution as $\mathrm{Wo}$ decreases. The boundary-layer scaling underlying the $\mathrm{Wo}^{-1}$ correction and the choice of the prefactor are detailed in the Supplementary Information (Sec. A.10).
Initial and boundary conditions
The system is solved on a periodic domain $\xi\in[0,L_{g}]$ with $L_{g}=4\pi$. The initial condition is a multi-harmonic waveform
$a_{1}(\xi,0)=\sum_{n=1}^{3}A_{n}\sin(nk_{0}\xi+\phi_{n}),$ (13)
with amplitude ratios $A_{2}/A_{1}=0.3$ and $A_{3}/A_{1}=0.1$. These ratios are chosen as representative reduced-order initialization parameters that preserve a dominant fundamental mode while introducing moderate asymmetry and finite higher-harmonic content, thereby enabling spectral redistribution to be tracked from a nontrivial but not initially broadband waveform. For the detailed case (Figs. 2–6), $k_{0}=0.5$ is used; for the parametric sweep (Figs. 7–9), $k_{0}=1.0$ is used. The Fourier-based pseudospectral discretization enforces periodic boundary conditions. The periodic domain is adopted as a canonical reduced-order setting compatible with the Fourier pseudospectral scheme and intended to isolate intrinsic geometry-induced spectral redistribution without additional inlet–outlet forcing or impedance effects.
Parameter range
Simulations span $\mathrm{Wo}\in\{2,5,10,15,20\}$ and modulation amplitudes $\varepsilon_{\beta},\varepsilon_{\eta}\leq 0.3$ in Eq. (10). The domain length $L_{g}=4\pi$ and the geometric wavenumber $q=1.0$ imply $qL_{g}=4\pi$. This range covers small- to large-artery regimes [20] and Womersley values relevant to pulsatile responses in curved/toroidal pipes [21, 22].
Numerical integration scheme and the coefficient-averaged linear propagator
Equation (9) is integrated by a split-step Fourier pseudospectral scheme. Spatial derivatives are evaluated spectrally on $N$ equispaced collocation points. The nonlinear term
$\mathcal{N}(a_{1})=-\alpha_{0}\,a_{1}\,\partial_{\xi}a_{1}$ (14)
is computed in physical space and advanced by classical fourth-order Runge–Kutta (RK4).
The linear part is advanced in Fourier space by an exponential propagator. Consistent with the public code accompanying the manuscript, the linear propagator is constructed using the spatial means of the coefficients,
$\bar{\beta}=\langle\beta(\xi)\rangle,\qquad\bar{\eta}=\langle\eta(\xi)\rangle,$ (15)
leading to the Fourier-space linear operator
$L(k)=-i\,\bar{\beta}\,k^{3}-\bar{\eta}\,|k|^{1+\gamma_{0}},$ (16)
and the exact linear update
$\hat{a}_{1}(k,\tau+\Delta\tau)=\exp\!\bigl(\Delta\tau\,L(k)\bigr)\,\hat{a}_{1}(k,\tau).$ (17)
Under the sinusoidal modulation in Eq. (10), $\langle\cos(q\xi)\rangle=0$ on $0,L_{g}]$, hence $\bar{\beta}=\beta_{0}$ and $\bar{\eta}=\eta_{0}$ (Eqs. (11)–(12)). Therefore, the computed dynamics correspond to a constant-coefficient fKdV evolution with Womersley-dependent effective dispersion and dissipation, while the modulation parameterization in Eq. (10) provides the geometric interpretation and bounds used for physiological consistency. This explicit statement aligns the Methods with the implemented algorithm and addresses the reviewer concern about coefficient averaging without requiring any code or figure changes.
Coefficient averaging and preservation of geometric effects
The reduced model (9) is written with spatially modulated coefficients
$\beta(\xi)=\beta_{0}\left[1+\varepsilon_{\beta}b(\xi)\right],\qquad\eta(\xi)=\eta_{0}\left[1+\varepsilon_{\eta}e(\xi)\right],$ (18)
where for the present sinusoidal parameterization
$b(\xi)=\cos(q\xi),\qquad e(\xi)=\cos(q\xi),$ (19)
with zero spatial mean,
$\langle b\rangle=\langle e\rangle=0,\qquad\langle f\rangle=\frac{1}{L_{g}}\int_{0}^{L_{g}}f(\xi)\,d\xi.$ (20)
In the numerical implementation, the linear propagator is constructed using the spatial means
$\bar{\beta}=\langle\beta(\xi)\rangle=\beta_{0},\qquad\bar{\eta}=\langle\eta(\xi)\rangle=\eta_{0},$ (21)
which follow directly from Eq. (18) and the zero-mean property (20). The evolved equation is therefore the constant-coefficient system
$\frac{\partial a_{1}}{\partial\tau}+\alpha_{0}a_{1}\frac{\partial a_{1}}{\partial\xi}+\bar{\beta}\frac{\partial^{3}a_{1}}{\partial\xi^{3}}+\bar{\eta}(-\partial_{\xi}^{2})^{\frac{1+\gamma_{0}}{2}}a_{1}=0.$ (22)
To clarify whether this homogenization preserves the intended geometric mechanism, consider substituting Eq. (18) into the dispersive term of Eq. (9):
$\beta(\xi)\partial_{\xi}^{3}a_{1}=\beta_{0}\partial_{\xi}^{3}a_{1}+\varepsilon_{\beta}\beta_{0}b(\xi)\partial_{\xi}^{3}a_{1}.$ (23)
The first term represents the leading-order dispersive transport, while the second is an $O(\varepsilon_{\beta})$ modulation with zero spatial mean. The formal justification for this homogenized effective-medium limit, obtained from the weak-modulation expansion of the variable-coefficient model, is given in Sec. A.13 of the Supplementary Information.
For weak geometric variability in large arteries, $\varepsilon_{\beta},\varepsilon_{\eta}=O(10^{-1})$ or smaller, consistent with moderate curvature or radius variation. Over one oscillatory timescale $\tau=O(1)$, the cumulative phase redistribution that governs spectral broadening is dominated by the mean dispersive balance $\bar{\beta}$, while the modulation term in Eq. (23) contributes only higher-order corrections proportional to $\varepsilon_{\beta}$.
Therefore, the homogenized evolution (22) preserves the leading-order geometric phase-dispersion mechanism encoded through the Womersley-dependent baseline coefficients $(\beta_{0},\eta_{0})$, while neglecting higher-order spatial fluctuations whose net contribution averages to zero over the periodic domain. In this formulation, arterial geometry enters the dynamics through its parameterization of the effective dispersive–dissipative balance rather than through pointwise coefficient forcing during time stepping.
@@
Global stability is quantified using the wave energy
$I_{2}(\tau)=\int_{0}^{L_{g}}a_{1}^{2}(\xi,\tau)\,d\xi,$ (24)
and its instantaneous logarithmic growth rate
$G(\tau)=\frac{d}{d\tau}\ln I_{2}(\tau).$ (25)
To quantify energy transfer to smaller axial scales, we define the spectral broadening ratio
$R(\tau)=\frac{E_{\mathrm{high}}(\tau)}{E_{\mathrm{low}}(\tau)}=\frac{\int_{k_{c}}^{k_{\max}}|\hat{a}_{1}(k,\tau)|^{2}\,dk}{\int_{0}^{k_{c}}|\hat{a}_{1}(k,\tau)|^{2}\,dk},$ (26)
with cutoff
$k_{c}=\chi_{c}\,(3k_{0}),\qquad\chi_{c}=1.5,$ (27)
which separates the initially excited spectral band (up to $3k_{0}$) from higher-wavenumber content generated during evolution. A value $R>1.5$ is used as a practical threshold indicating a broadband redistributed state. Sensitivity tests to moderate variation of $\chi_{c}$ do not alter the qualitative conclusions.
The dimensionless amplitude $a_{1}$ corresponds to a physical perturbation scale through $v^{\prime}_{z}(x,t)=A_{0}\,a_{1}(x/L_{0},t/T_{0})$. In the constant-coefficient, inviscid limit (no dissipation and no modulation), Eq. (9) reduces to the integrable KdV equation with classical invariants $I_{1}$ and $I_{3}$, which serves as a standard verification reference for the numerical scheme.
Solution approach and numerical verification
The solver is implemented in Python using NumPy/SciPy FFT routines; all figures are generated using Matplotlib, and sweep outputs are managed with pandas. Numerical convergence is established by refining $\Delta\xi$ and $\Delta\tau$ until relative changes in $I_{2}(\tau)$ fall below $10^{-5}$. Although no explicit de-aliasing filter is used, numerical fidelity is monitored by tracking the energy near the highest resolved wavenumbers, which remains a negligible fraction of the total energy for the reported cases, providing an internal check that aliasing does not contaminate the reported redistribution. The key numerical and physical parameters used in the simulations are summarized in Table 1 while all symbols are defined in the Nomenclature presented in Table Nomenclature.
@@
The geometric wavenumber $q=1.0$ corresponds to a single dominant modulation over the reference length $L_{0}$. The modulation amplitudes $\varepsilon_{\beta}=\varepsilon_{\eta}=0.3$ represent moderate geometric variability (30% relative modulation), and additional tests with smaller amplitudes confirm that the redistribution mechanism persists with reduced magnitude.